Moira is three years less than half her mother's age. Her mother is 76 years old. How old is Moira?
Answer by LinnW(1048)

set d = Moira the daughter's age
m = mother's age
d = -3 + (1/2)*m
m = 76
substitute 76 for m in d = -3 + (1/2)*m
d = -3 + (1/2)*76
d = -3 + 38
d = 35